# LeanKAN: A Parameter-Lean Kolmogorov-Arnold Network Layer with Improved Memory Efficiency and Convergence Behavior

### v1

@@ Abstract
The recently proposed Kolmogorov-Arnold network (KAN) is a promising alternative to multi-layer perceptrons (MLPs) for data-driven modeling. While original KAN layers were only capable of representing the addition operator, the recently-proposed MultKAN layer combines addition and multiplication subnodes in an effort to improve representation performance. Here, we find that MultKAN layers suffer from a few key drawbacks including limited applicability in output layers, bulky parameterizations with extraneous activations, and the inclusion of complex hyperparameters. To address these issues, we propose LeanKANs, a direct and modular replacement for MultKAN and traditional AddKAN layers. LeanKANs address these three drawbacks of MultKAN through general applicability as output layers, significantly reduced parameter counts for a given network structure, and a smaller set of hyperparameters. As a one-to-one layer replacement for standard AddKAN and MultKAN layers, LeanKAN is able to provide these benefits to traditional KAN learning problems as well as augmented KAN structures in which it serves as the backbone, such as KAN Ordinary Differential Equations (KAN-ODEs) or Deep Operator KANs (DeepOKAN). We demonstrate LeanKAN’s simplicity and efficiency in a series of demonstrations carried out across both a standard KAN toy problem and a KAN-ODE dynamical system modeling problem, where we find that its sparser parameterization and compact structure serve to increase its expressivity and learning capability, leading it to outperform similar and even much larger MultKANs in various tasks.
I Introduction
@@
Kolmogorov-Arnold Networks (KANs) [1, 2] are an emerging tool for data-driven and physics-inspired modeling. Traditional multilayer perceptrons (MLPs) contain predefined activation functions and learnable weights and biases to learn function approximations. KANs, on the other hand, apply the Kolmogorov-Arnold Theorem [3, 4] to instead learn the activation functions themselves, through gridded sums of univariate basis functions. These unique univariate activations are learned across all inputs, are then combined via parameter-free summation at each output. In other words, MLPs learn weights and biases with fixed activation functions, while KANs learn activation functions with fixed weights and biases. See Fig. 1(A) for a visual representation. This characteristic of learnable activation functions in KANs has shown to provide increased interpretability thanks to their visualizable profiles, rapid convergence with small parameter sizes thanks to their improved neural scaling laws, and the potential for human-readable output via sparse symbolic regression [1, 2, 5]. Mirroring the development trajectory of MLP-based networks through various augmented frameworks, KANs have been applied in architectures such as physics informed KANs [6, 7, 8], convolutional KANs [9], Deep Operator KANs [10], and KAN Ordinary Differential Equations [5]. Such applications retain the key inductive biases and physical insights embedded in these specialized structures (as originally studied in the standard MLP variants), while enabling the various inference benefits inherent to KANs, such as increased interpretability and high-order neural convergence.
While these applications have shown significant promise, the original KAN formulation of Liu et al. [1] on which they were built contains a key limitation to expressivity. With learnable univariate activations combined at each output only via addition, it is straightforward to express complex sums of univariate functions on two inputs $x$ and $y$, such as $e^{sin(x^{2}+y^{2})}$, through gridded univariate basis functions combined via addition. It is surprisingly challenging, however, to represent even trivial multiplication such as $xy$. A theoretical two-input, one-output KAN layer combines the inputs $x$ and $y$ via learnable activations $f_{a}(x)$ and $g_{a}(y)$, to construct an output of the form $f_{a}(x)+g_{a}(y)$. To represent $xy$, there does not exist a single-layer KAN solution due to the lack of multiplication expression across inputs. To get around this, a two-layer KAN was seen in [1] to converge to the roundabout $xy=((x+y)^{2}-(x^{2}+y^{2}))/2$ using four activations in the first layer, which is needlessly bulky especially in light of the typically sparse and shallow exact function reconstructions presented in the remainder of that work. This naturally called for the follow-up work (KAN 2.0) of Liu et al. [2], where MultKAN was proposed. This augmented structure begins with a similar formulation to a standard KAN layer (henceforth referred to as AddKAN), but adds additional activation functions to enable an additional multiplication sublayer. This sublayer directly encodes the multiplication operator, allowing for simple and easily interpretable expression of trivial multiplication-based functions, like $xy$, up to more complex functions involving multiplication, like $sin(x)y^{2}$, with a single layer only. More mathematical detail is provided in Sec. II.2, with a visualization in Fig. 1(B). In the same example where AddKAN learns the activation functions $f_{a}(x)$ and $g_{a}(y)$ for an output value $f_{a}(x)+g_{a}(y)$, a single MultKAN layer would additionally learn the activation functions $f_{m}(x)$ and $g_{m}(y)$ in parallel, combine them via $f_{m}(x)+g_{m}(y)$, and then inject the multiplication operator by computing the final output as $[f_{a}(x)+g_{a}(y)]\cdot[f_{m}(x)+g_{m}(y)]$, i.e. the product of the original AddKAN output node with a newly-added MultKAN output node. Here, we see that $xy$ can be learned trivially in a single layer using $f_{a}(x)=x,g_{a}(y)=0,f_{m}(x)=0,$ and $g_{m}(y)=y$. To summarize, where AddKAN requires two layers and six activation functions to learn a roundabout form of simple multiplication between two inputs, MultKAN does so with a single layer comprising four activations (two of which are trivially equal to zero). The improvements in expressivity and memory efficiency resulting from this formulation change are evident. Additionally, the substantial reduction in bulk and complexity offers notable advantages in symbolic regression and sparse representation, which are crucial for many KAN applications [1, 2, 5].
Despite the partial success of MultKAN, its specific construction bars it from general use in multidimensional outputs (i.e., $[x^{3}y,x+y,xy^{2}]$ rather than simply $xy$). Certain output nodes see the multiplication operator while others are restricted to the addition operator only, which we demonstrate in this work to be a significant training bottleneck. For example, the hyperparameters in a MultKAN network may dictate that only the second and third output values are learned using the multiplication operator, making the second and third indices of $[x^{3}y,x+y,xy^{2}]$ easily learnable in a single layer but the first index impossible to learn for the same reason that AddKAN cannot learn $xy$. This was not an issue in the original MultKAN manuscript [2] where investigation involved scalar outputs only, but constrains the applicability and interpretability of MultKAN in many real physical problems of KAN modeling interest with multidimensional outputs [5, 11]. In a similar vein, the MultKAN implementation adds two new hyperparameters, increasing the complexity of potential user tuning with this new framework. Finally, in terms of quantitative performance drawbacks, the architecture of a MultKAN dictates that for a given node and grid structure, additional activation functions are required beyond what is needed for AddKAN, degrading the memory efficiency of the MultKAN structure and potentially impacting its training speed and convergence behavior. In the example discussed above, for example, the two-input, one-output AddKAN learns only two activations, $f_{a}(x)$ and $g_{a}(y)$, while the two-input, one-output MultKAN learns $f_{a}(x),f_{m}(x),g_{a}(y),g_{m}(y)$. It is clear that the MultKAN performs better in this case thanks to the multiplication operator enabling it to learn $xy$, but it also “wastes” two activations to do so: $g_{a}(y)=0,f_{m}(x)=0$. In all KAN layer structures, the number of parameters scales proportionally with the number of activations (while the addition and multiplication of activations are parameter-free), so the MultKAN here would appear to have twice the number of parameters necessary in the obvious case of learning $xy$ via $f_{a}(x)\cdot g_{a}(y)$, with $f$ and $g$ as identity functions.
In this work, we propose LeanKAN which substantially improves expressivity, memory efficiency, and training speeds of the network. This entirely reformulated multiplication structure for KAN layers allows us to eliminate one of two added hyperparameters, completely revert the increase in parameter size (returning to a size equal to that of the corresponding AddKAN), and provide general use in KANs with multidimensional outputs while retaining the multiplication operator. An example LeanKAN structure is shown in Fig. 1(C). LeanKAN eliminates the $f_{m}(x)$ and $g_{m}(y)$ activations, which we find here to be unnecessary for effective training, allowing for direct learning of $xy=f_{a}(x)\cdot g_{a}(y)$. As a replacement for individual KAN layers themselves, we expect that LeanKAN can be widely adapted in currently active fields of PDE discovery and surrogate modeling without restriction to specific models. Therefore, any performance benefits have potential impact not only for improved efficiency in standard KAN applications as a function approximator, but also in improved efficiency as the backbone of any number of these augmented KAN structures such as physics-informed KAN, DeepOKAN, and KAN-ODEs.
@@
Accordingly, we compare LeanKAN against the original MultKAN in a series of tests within two case studies: one in a standard KAN learning application to demonstrate the specific structural benefits of LeanKAN, and the other in a KAN-ODE application to demonstrate how these benefits can scale to already-developed, augmented KAN structures. In all cases, we find that the reduced activation and parameter size of LeanKAN does not deteriorate training dynamics or converged performance. In fact, the more compact and parameter-lean LeanKAN networks achieve slightly stronger metrics across the board, thanks to their more efficient use of parameters and activations.
II Background
@@
Figure 1: Visualization of the three KAN layers referenced in this work, all with four inputs and three outputs. (A) A standard AddKAN [1]. (B) A standard MultKAN [2] with second-order multiplication ($k=2$), and one addition/identity output node ($n^{a}=1$). (C) Currently proposed LeanKAN with two multiplication input nodes ($n^{mu}=2$). Black nodes indicate inputs, while all others are labeled with addition, multiplication, or blank identity operators. Note that all KAN network parameters are defined at the activation functions only, and all connections and operators are parameter-free. Opacity of activations and connections is varied to imply the relative importance of different activations (as in [1]). Note that MultKAN requires 20 activations compared to AddKAN and LeanKAN’s 12 (with a proportional increase in parameters and memory), with even more required for larger $n^{a}$ or $k$ hyperparameters.
@@
Number of Activation Functions ($\propto$ Number of Parameters) $n_{l}\cdot n_{l+1}$ $n_{l}\cdot n_{l+1}+\underbrace{n_{l}(k-1)n^{m}}_{increased}$ $n_{l}\cdot n_{l+1}$
@@
Multiplication N/A $n^{a}$ (# of additive nodes) $n^{mu}$ (# of multiplicative nodes
@@
Reference Ref. [1] Ref. [2] This work
II.1 Traditional Kolmogorov-Arnold Network (AddKAN or KAN 1.0)
Here, $\phi_{q,p}:~{}[0,1]\rightarrow\mathbb{R}$ are the univariate activation functions evaluated on the inputs $x_{p}$ to construct the $q^{th}$ node of the hidden layer, while $\Phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$ are the second layer’s activation functions that construct the scalar output from the hidden layer node values (overall representing a two-layer KAN with a hidden layer width of $12n+$ 1 and an output dimension of 1). In matrix form, the $n_{l+1}$-dimensional output of a single KAN layer with $n_{l}$ inputs can be represented as
$\mathbf{y}=\Phi_{l}\mathbf{x}\in\mathbb{R}^{n_{l+1}},$ (2)
where $\phi$ is the activation function with subscripts $l$ indicating the index of the current layer, $n_{l}$ indicating the number of nodes in the current layer, and $n_{l+1}$ indicating the number of nodes in the subsequent layer. In other words, each input is connected to each output with a unique learnable activation function, leading to a total of $n_{l}\cdot n_{l+1}$ activation functions connecting the $l^{\text{th}}$ and $(l+1)^{\text{th}}$ layers. An AddKAN layer with four inputs and three outputs is shown in Fig. 1(A), with a corresponding 12 activation functions. Finally, by stacking this AddKAN layer (Eq. 2), a multi-layer KAN can be constructed such that $\text{KAN}(\mathbf{x})=\left({\Phi}_{L-1}\circ{\Phi}_{L-2}\circ\cdots\circ{\Phi}_{1}\circ{\Phi}_{0}\right)\mathbf{x}$.
The original KAN implementation used B-splines as the basis functions that comprise each activation, although it was later shown in Li [13] that Gaussian radial basis functions (RBFs) are more efficient. Here, we define the individual activation functions as
$\phi_{l,\alpha,\beta}\left(\text{x}\right)$ $=\sum_{i=1}^{N}w^{\psi}_{l,\alpha,\beta,i}\cdot\psi\left(\lvert\lvert\text{x}-c_{i}\rvert\rvert\right)+w^{b}_{l,\alpha,\beta}\cdot b\left(\text{x}\right),$ (4)
$\psi(r)$ $=\exp(-\frac{r^{2}}{2h^{2}}),$ (5)
where $N$ is the grid size (number of gridded basis functions used to reconstruct a single activation), $w^{\psi}_{l,\alpha,\beta,i}$ and $w^{b}_{l,\alpha,\beta}$ are the learnable parameters that scale and superimpose the RBF basis functions $\psi(r)$ and the base activation function $b(\text{x})$, respectively, $\alpha$ and $\beta$ are the index of the layer $l$’s matrix in Eq. 3 on which these weights are applied, $c_{i}$ are the centerpoints of each gridded basis function, and $h$ is the gridpoint spacing (i.e. the RBF spreading parameter). Normalization is carried out at each layer as in Refs. [14, 15], to avoid the original and expensive re-gridding technique of Liu et al. [1]. Base activations $b(\text{x})$ are directly defined as Swish activation functions [16].
II.2 Multiplication Kolmogorov-Arnold Network (MultKAN or KAN 2.0)
The subsequent work of the original KAN authors [1] introduced MultKAN [2], which augments the summation terms in Eq. 1 with optional multiplication connections, allowing for more inherent expressivity while training and when symbolically expressing functions. MultKAN starts from an identical construction as AddKAN, where an input $\mathbf{x}$ is transformed via the activation function matrix $\Phi$ into $\mathbf{y}$ (as in Eqs. 2 and 3), where $\mathbf{y}$ is defined here as a hidden subnode output rather than the final layer output. Then, to compute the true layer output $\mathbf{z}$, an auxiliary activation-free (and thus parameter-free) layer $\Omega_{l}$ is defined in Eq. 6 as per the following notation in Python for $l^{\text{th}}$ layer.
$\mathbf{z}_{l}=\mathbf{\Omega_{l}}\mathbf{y}_{l}=[\underbrace{\mathbf{y}_{l}[1:n^{a}_{l+1}]}_{\text{identity}},\underbrace{{\mathbf{y}_{l}[n^{a}_{l+1}::2]\odot\mathbf{y}_{l}[n^{a}_{l+1}+1::2]}}_{\text{MultKAN segment ({\hbox{k=2}})}}]\in\mathbb{R}^{n^{a}_{l+1}+n^{m}_{l+1}}.$ (6)
Here, the sublayer $\mathbf{y}_{l}$ with $n_{l+1}$ output nodes is split into $n_{l+1}^{a}$ addition-only nodes and $n_{l+1}^{m}=n_{l+1}-n_{l+1}^{a}$ multiplication nodes. The mapping from $\mathbf{y}$ to $\mathbf{z}$ performs no operations for the first $n_{l+1}^{a}$ identity nodes of $\mathbf{z}$ (“identity”). For the remaining $n_{l+1}^{m}$ nodes of $\mathbf{z}$, however, $k^{\text{th}}$ order multiplication is carried out on the subnode $y$ values (“MultKAN segment”), where we define “order” in this case as the number of activations included in the product (i.e. $y_{1}\cdot y_{2}$ is second order, $y_{1}\cdot y_{2}\cdot y_{3}$ is third order). Here the hyperparameter $k=2$ leads to second order, pairwise multiplication as denoted by $\odot$, while higher values of $k$ would encode higher-order multiplication terms in the MultKAN segment of this equation. A MultKAN layer with four inputs and three outputs is shown in Fig. 1(B), where $n_{l+1}^{a}$=1 dicates that $z_{1}=y_{1}$ (“identity”), while $z_{2}$ and $z_{3}$ are the respective pairwise products of $y_{2}$ and $y_{3}$; and $y_{4}$ and $y_{5}$ (“MultKAN segment”). To facilitate pairwise multiplication for $z_{2}$ and $z_{3}$ with this set of hyperparameters, we see that two additional nodes in the $\mathbf{y}$ layer are required. These added nodes increase the number of activations to 20, a $66\%$ increase from the 12 needed in the same structure for AddKAN (see Fig. 1(A)). We note that in any case with $n_{l+1}^{m}\neq 0$, the dimension of $\mathbf{y}\in\mathbb{R}^{n^{a}_{l+1}+kn^{m}_{l+1}}$ is larger than the dimension of $\mathbf{z}\in\mathbb{R}^{n^{a}_{l+1}+n^{m}_{l+1}}$ by $(k-1)n^{m}_{l+1}$, in order to facilitate these multiplication connections. This leads to a proportional increase in activation functions (and identically in the number of parameters).
Due to the increased number of activations needed for multiplication, MultKAN exhibits inflation of the total number of parameters in a given layer, especially for higher-order multiplication (i.e. $k>2$). In other words, MultKAN has a parameter count strictly larger than a standard AddKAN, even for an identical layer structure (input nodes, output nodes, and grid size). Therefore, while it benefits from the inclusion of the multiplication operator and any inherent representation capability or expressivity contained therein, the MultKAN structure is also hypothesized here to deteriorate memory efficiency and reduce training speeds as a result of this parameter inflation.
Another limitation of this MultKAN implementation is that it is not generally applicable in the last layer of a multi-output KAN. We see in Eq. 6 that the array of output nodes $\mathbf{z}_{l}$ has certain indices (up to $n^{a}_{l+1}$) that include only identity transformations of the subnode inputs $\mathbf{y}_{l}$, while other indices ($n^{a}_{l+1}+1$ to the end) include multiplicative transformations of the subnode inputs. This asymmetrical structure can lead to performance degradation with multidimensional output layers, where for example if an output vector includes $z_{i}=x_{1}\cdot x_{2}$ where $i<n^{a}$, i.e. $z_{i}=y_{i}$ with no multiplication terms, then MultKAN has effectively reverted to AddKAN due to the lack of multiplication on the $\mathbf{y}$ sublayer for the $z_{i}$ output, making $x_{1}\cdot x_{2}$ impossible to learn in a single layer.
III Method
III.1 Parameter-Lean Kolmogorov-Arnold Network (LeanKAN)

### v2

@@ Abstract
The recently proposed Kolmogorov-Arnold network (KAN) is a promising alternative to multi-layer perceptrons (MLPs) for data-driven modeling. While original KAN layers were only capable of representing the addition operator, the recently-proposed MultKAN layer combines addition and multiplication subnodes in an effort to improve representation performance. Here, we find that MultKAN layers suffer from a few key drawbacks including limited applicability in output layers, bulky parameterizations with extraneous activations, and the inclusion of complex hyperparameters. To address these issues, we propose LeanKANs, a direct and modular replacement for MultKAN and traditional AddKAN layers. LeanKANs address these three drawbacks of MultKAN through general applicability as output layers, significantly reduced parameter counts for a given network structure, and a smaller set of hyperparameters. As a one-to-one layer replacement for standard AddKAN and MultKAN layers, LeanKAN is able to provide these benefits to traditional KAN learning problems as well as augmented KAN structures in which it serves as the backbone, such as KAN Ordinary Differential Equations (KAN-ODEs) or Deep Operator KANs (DeepOKAN). We demonstrate LeanKAN’s simplicity and efficiency in a series of demonstrations carried out across a standard KAN toy problem as well as ordinary and partial differential equations learned via KAN-ODEs, where we find that its sparser parameterization and compact structure serve to increase its expressivity and learning capability, leading it to outperform similar and even much larger MultKANs in various tasks.
keywords:
Kolmogorov-Arnold Networks , Machine Learning , Interpretable Networks , Model Discovery , Data-Driven Modeling
1 Introduction
Kolmogorov-Arnold Networks (KANs) are an emerging tool for data-driven and physics-inspired modeling. The Kolmogorov-Arnold representation theorem (KAT) [14, 11], as a central idea of KANs, states that any multivariate function can be represented by the superposition of continuous univariate functions. Early work proposed the use of cubic spline basis functions as the building blocks for KAT compositions [10, 5]. Recent works have proposed combining the KAT with descent methods as a multi-layer deep learning approach, optimizing piecewise linear [19, 20], ReLu [26], and other arbitrary splines [21] to approximate desired solutions. Following these pioneering works, Liu et al. demonstrated a parallel capability of KAT-based representations in their human-interpretable learned activations facilitated by pruning, sparsification, and symbolic regression [17].
@@ 1 Introduction
KANs have received increasing attention recently thanks to favorable qualities compared to standard MLP networks. They provide rapid convergence with small parameter sizes thanks to their improved neural scaling laws, interpretability thanks to their visualizable profiles, and the potential for human-readable output via sparse symbolic regression [17, 16, 13]. Mirroring the development trajectory of MLP-based networks through various augmented frameworks, KANs have been applied in architectures such as physics informed KANs [18, 7, 9], convolutional KANs [3], Deep Operator KANs [1], and KAN Ordinary Differential Equations [13]. Such applications retain the key inductive biases and physical insights embedded in these specialized structures (as originally studied in the standard MLP variants), while enabling the various inference benefits inherent to KANs, such as increased interpretability and high-order neural convergence.
While these applications have shown significant promise, the original KAN formulation of Liu et al. [17] (henceforth denoted AddKAN in Fig. 1(A)) on which they were built contains a key limitation to expressivity due to its use of summation only to combine the learnable activation functions. With this framework it is straightforward to express additive functions on two inputs $x$ and $y$ such as $e^{\sin(x^{2}+y^{2})}$, but surprisingly challenging to express the simple product $xy$. In fact, the latter is impossible to learn with a single addition layer, while a two-layer network was shown in [17] to require six activations to learn the roundabout and bulky $xy=((x+y)^{2}-(x^{2}+y^{2}))/2$. This naturally called for the follow-up work (KAN 2.0) of Liu et al. [16], where MultKAN was proposed. To directly encode the multiplication operator, MultKAN injects extra addition-only activations to create a summed sublayer ouput of higher dimension than the true output, then combines these sublayers through a hyperparameter-defined split of identity and multiplication operations to arrive at the true output. With MultKAN, $xy$ (and other arbitrary products) can be represented with a single layer only. More mathematical detail is provided in Sec. 2.2, with a visualization in Fig. 1(B). Notably, however, while MultKAN eliminates a layer when compared to the AddKAN implementation, its augmentation with increased activations and an additional sublayer appears to introduce extraneous parameters and learning bulk beyond what is necessary to efficiently express many learning problems. Additionally, the full definition of the new sublayer requires the addition of two hyperparameters not used in the standard AddKAN. Finally, while originally studied only for scalar solution quantities, the multidimensional generalization of MultKAN contains restricted multiplicative expressivity due to its separate mathematical treatment of the nodes augmented with multiplication and the remaining nodes left with addition and identity operations only.
@@
Accordingly, we compare LeanKAN against the original MultKAN across three case studies: a standard KAN learning application to demonstrate the specific structural benefits of LeanKAN; a detailed analysis in a KAN-ODE application to demonstrate how these benefits can scale to already-developed, augmented KAN structures; and a final PDE study leveraging KAN-ODEs to demonstrate LeanKAN’s scalability to larger problems. In all cases, we find that the reduced activation and parameter size of LeanKAN does not deteriorate training dynamics or converged performance. In fact, the more compact and parameter-lean LeanKAN networks achieve slightly stronger metrics across the board, thanks to their more efficient use of parameters and activations.
2 Background
@@
Figure 1: Visualization of the three KAN layers referenced in this work, all with four inputs and three outputs. (A) A standard AddKAN [17] with outputs $\mathbf{y}$. (B) A standard MultKAN [16] using second-order multiplication ($k=2$) and one addition/identity output node ($n^{a}=1$), with outputs $\mathbf{z}$. (C) Currently proposed LeanKAN using two multiplication input nodes ($n^{\text{mu}}=2$), with outputs $\mathbf{z}$. Black nodes indicate inputs, while all others are labeled with addition, multiplication, or blank identity operators. Note that all KAN network parameters are defined at the activation functions only, and all connections and operators are parameter-free. Opacity of activations and connections is varied to imply the relative importance of different activations (as in [17]). Note that MultKAN requires 20 activations compared to AddKAN and LeanKAN’s 12 (with a proportional increase in parameters and memory), with even more required for smaller $n^{a}$ or larger $k$ hyperparameters.
2.1 Traditional Kolmogorov-Arnold Network (AddKAN or KAN 1.0)
Here, $\phi_{q,p}:~{}[0,1]\rightarrow\mathbb{R}$ are the univariate activation functions evaluated on the inputs $x_{p}$ to construct the $q^{\text{th}}$ node of the hidden layer, while $\Phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$ are the second layer’s activation functions that construct the scalar output from the hidden layer node values (overall representing a two-layer KAN with a hidden layer width of $12n+$ 1 and an output dimension of 1). In matrix form, the $n_{l+1}$-dimensional output of a single KAN layer with $n_{l}$ inputs can be represented as
$\mathbf{y}=\Phi_{l}\mathbf{x}\in\mathbb{R}^{n_{l+1}},$ (2)
where $\phi$ is the activation function with subscripts $l$ indicating the index of the current layer, $n_{l}$ indicating the number of nodes in the current layer, and $n_{l+1}$ indicating the number of nodes in the subsequent layer. This formulation encodes unique connections between each distinct input-output pair, leading to a total of $n_{l}n_{l+1}$ activation functions connecting the $l^{\text{th}}$ and $(l+1)^{\text{th}}$ layers. An AddKAN layer with four inputs and three outputs is shown in Fig. 1(A), with 12 corresponding activation functions. Finally, by stacking this AddKAN layer (Eq. 2), a multi-layer KAN can be constructed such that $\text{KAN}(\mathbf{x})=\left({\Phi}_{L-1}\circ{\Phi}_{L-2}\circ\cdots\circ{\Phi}_{1}\circ{\Phi}_{0}\right)\mathbf{x}$.
The original KAN implementation used B-splines as the basis functions that comprise each activation, although it was later shown [15] that Gaussian radial basis functions (RBFs) are more efficient. Here, we define the individual activation functions as
$\phi_{l,\alpha,\beta}\left(\text{x}\right)$ $=\sum_{i=1}^{N}w^{\psi}_{l,\alpha,\beta,i}\psi\left(\lvert\lvert\text{x}-c_{i}\rvert\rvert\right)+w^{b}_{l,\alpha,\beta}b\left(\text{x}\right),$ (4)
$\psi(r)$ $=\exp(-\frac{r^{2}}{2h^{2}}),$ (5)
where $N$ is the grid size (number of gridded basis functions used to reconstruct a single activation), $w^{\psi}_{l,\alpha,\beta,i}$ and $w^{b}_{l,\alpha,\beta}$ are the learnable parameters that scale and superimpose the RBF basis functions $\psi(r)$ and the base activation function $b(\text{x})$, respectively, $\alpha$ and $\beta$ are the index of the layer $l$’s matrix in Eq. 3 on which these weights are applied, $c_{i}$ are the centerpoints of each gridded basis function, and $h$ is the gridpoint spacing (i.e. the RBF spreading parameter). Normalization is carried out at each layer as in prior works [2, 22], to avoid the original and expensive re-gridding technique of Liu et al. [17]. Base activations $b(\text{x})$ are directly defined as Swish activation functions [25].
2.2 Multiplication Kolmogorov-Arnold Network (MultKAN or KAN 2.0)
Subsequent work introduced MultKAN [16], which augments the summation terms in Eq. 1 with optional multiplication connections, allowing for more inherent expressivity while training and when symbolically expressing functions. MultKAN starts from an identical construction as AddKAN, where an input $\mathbf{x}$ is transformed via the activation function matrix $\Phi$ into $\mathbf{y}$ (as in Eqs. 2 and 3), which is defined here as a hidden subnode output rather than the final layer output. Then, to compute the true layer output $\mathbf{z}$, an auxiliary activation-free (and thus parameter-free) layer $\Omega_{l}$ is defined in Eq. 6 as per the following notation in Python notation for $l^{\text{th}}$ layer.
$\mathbf{z}_{l}=\mathbf{\Omega_{l}}\mathbf{y}_{l}=[\underbrace{\mathbf{y}_{l}[1:n^{a}_{l+1}]}_{\text{identity}},\underbrace{{\mathbf{y}_{l}[n^{a}_{l+1}::2]\odot\mathbf{y}_{l}[n^{a}_{l+1}+1::2]}}_{\text{MultKAN segment ($k=2$)}}]\in\mathbb{R}^{n^{a}_{l+1}+n^{m}_{l+1}}.$ (6)
An identical formulation can be written in standard mathematical notation as,
${{z}_{l,i}}={{y}_{l,i}}$ $\text{for}~{}i\in\{1,2,...,n^{a}_{l+1}\}\subset\mathbb{N},$ (7)
${{z}_{l,i}}={{y}_{l,2i-n^{a}_{l+1}-1}{y}_{l,2i-n^{a}_{l+1}}}$ $\text{for}~{}i\in\{n^{a}_{l+1}+1,n^{a}_{l+1}+2,...,n^{a}_{l+1}+n^{m}_{l+1}\}\subset\mathbb{N}.$ (8)
In Eqs. 6-8, the output $\mathbf{z}_{l}$ with $n_{l+1}$ nodes is split into $n_{l+1}^{a}$ addition-only nodes and $n_{l+1}^{m}=n_{l+1}-n_{l+1}^{a}$ multiplication nodes. The mapping from $\mathbf{y}$ to $\mathbf{z}$ performs no operations for the first $n_{l+1}^{a}$ identity nodes of $\mathbf{z}$ (“identity” in Eq. 6; Eq. 7). For the remaining $n_{l+1}^{m}$ nodes of $\mathbf{z}$, however, $k^{\text{th}}$ order multiplication is carried out on the subnode $y$ values (“MultKAN segment” in Eq. 6; Eq. 8), where we define “order” in this case as the number of activations included in the product (i.e. $y_{1}y_{2}$ is second order, $y_{1}y_{2}y_{3}$ is third order). Here the hyperparameter $k=2$ leads to second order, pairwise multiplication as denoted by $\odot$ in Eq. 6, while higher values of $k$ would encode higher-order multiplication terms in the MultKAN segment of this formulation (and lower values of $k$ are not possible). It is clear in Eq. 8 that $\mathbf{y}$ contains $n^{m}_{l+1}$ more nodes than $\mathbf{z}$ for this case with $k=2$, implying a proportionally increased number of activations and network parameters.
A MultKAN layer with four inputs and three outputs is shown in Fig. 1(B), where $n_{l+1}^{a}$=1 dicates that $z_{1}=y_{1}$ (“identity”), while $z_{2}$ and $z_{3}$ are the respective pairwise products of $y_{2}$ and $y_{3}$; and $y_{4}$ and $y_{5}$ (“MultKAN segment”). To facilitate pairwise multiplication for $z_{2}$ and $z_{3}$ with this set of hyperparameters, we see that two additional nodes in the $\mathbf{y}$ layer are required. These added nodes increase the number of activations to 20, a $66\%$ increase from the 12 needed in the same structure for AddKAN (see Fig. 1(A)). In fact, in all nontrivial cases ($n_{l+1}^{m}\neq 0$), the dimension of $\mathbf{y}\in\mathbb{R}^{n^{a}_{l+1}+kn^{m}_{l+1}}$ is larger than the dimension of $\mathbf{z}\in\mathbb{R}^{n^{a}_{l+1}+n^{m}_{l+1}}$ by $(k-1)n^{m}_{l+1}$, in order to facilitate these multiplication connections. The increased number of activations in the sublayer leads to a parameter count strictly larger than a standard AddKAN. This is true even for an identical layer structure (input nodes, output nodes, and grid size), and is exacerbated with higher-order multiplication (i.e. $k>2$). Therefore, while the MultKAN structure benefits from the inclusion of the multiplication operator and any inherent representation capability or expressivity contained therein, it is also hypothesized here to deteriorate memory efficiency and reduce training speeds as a result of this parameter inflation.
Another limitation of this MultKAN implementation is that it is not generally applicable in the last layer of a multi-output KAN. We see in Eq. 6 that the array of output nodes $\mathbf{z}_{l}$ has certain indices (up to $n^{a}_{l+1}$) that include only identity transformations of the subnode inputs $\mathbf{y}_{l}$, while other indices ($n^{a}_{l+1}+1$ to the end) include multiplicative transformations of the subnode inputs. This asymmetrical structure can lead to performance degradation with multidimensional output layers, where, for example, an output node $z_{i}=x_{1}x_{2}$ is impossible to learn in a single layer if $i\leq n^{a}$, as the MultKAN at that $i$ does not encode multiplication.
@@
Num. activation functions ($\propto$ num. parameters) $n_{l}n_{l+1}$ $n_{l}n_{l+1}+\underbrace{n_{l}(k-1)n^{m}}_{\text{increased}}$ $n_{l}n_{l+1}$
@@
Multiplication N/A $n^{a}$ (# of additive nodes) $n^{\text{mu}}$ (# of multiplicative nodes
@@
Reference [17] [16] Our work
3 Method
3.1 Parameter-Lean Kolmogorov-Arnold Network (LeanKAN)